Really need help with these. Can someone show me how to break these problems down and slove step by step. Thank you!
I THINK 1,2, AND 3 ARE FOIL METHOD, AM NOT SURE
1) 1X (2X+3)
2) 0X (2X+5)
3) 0:(2X+3)
4) 16:0
5) FIND THE SLOPE OF THE EQUATION: 4Y= 4X+4
6) fIND THE SLOPE OF THE EQUATION: X+Y=7

Answer by stanbon(75887)

1) 1X (2X+3) = 2x+3
--------------------
2) 0X (2X+5) = 0
--------------------
3) 0:(2X+3) = 0
--------------------
4) 16:0 is undefined
--------------------
5) FIND THE SLOPE OF THE EQUATION:
4Y= 4X+4
y = x + 1
slope = 1
------------------------
6) fIND THE SLOPE OF THE EQUATION: X+Y=7
y = -x + 7
slope = -1
=================================
Cheers,
Stan H.
==========